40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 19 22:05:09 2016, Koji, Summary, IOO, PMC, IMC aligned. The ringdown PD/Lens removed. 
    Reply  Wed Dec 21 21:03:06 2016, Koji, Summary, IOO, IMC WFS tuning DSC_0142.JPG
       Reply  Thu Dec 22 18:39:09 2016, Koji, Summary, IOO, IMC WFS tuning WFS_error_noise.pdfWFS_error_noise_chans.pdf
          Reply  Fri Dec 23 20:53:44 2016, Koji, Summary, IOO, IMC WFS tuning DSC_0144.JPGDSC_0145.JPG
             Reply  Fri Dec 23 21:05:56 2016, Koji, Summary, IOO, IMC WFS tuning IMC_WFS_segment_TF.pdfIMC_WFS_channels_TF.pdfIMC_WFS_161221_table1.pdfIMC_WFS_161221_table2.pdfIMC_WFS_161221.xlsx.zip
                Reply  Sun Dec 25 14:39:59 2016, Koji, Summary, IOO, IMC WFS tuning 
                   Reply  Mon Dec 26 12:45:31 2016, Koji, Summary, IOO, IMC WFS tuning IMC_REFL_TRANS_26hrs.pngIMC_TRANS_P_Y_26hrs.png
          Reply  Thu Dec 29 13:22:21 2016, rana, Summary, IOO, IMC WFS tuning 
             Reply  Thu Dec 29 16:52:51 2016, Koji, Summary, IOO, IMC WFS tuning 
                Reply  Thu Dec 29 21:35:30 2016, rana, Summary, IOO, IMC WFS tuning Screen_Shot_2017-01-03_at_5.55.13_PM.png
                   Reply  Tue Jan 24 01:04:16 2017, gautam, Summary, IOO, IMC WFS RF power levels MC1_misalignment.png
                      Reply  Fri Jan 27 00:14:02 2017, gautam, Summary, IOO, IMC WFS RF power levels 
Message ID: 12679     Entry time: Mon Dec 19 22:05:09 2016     Reply to this: 12680
Author: Koji 
Type: Summary 
Category: IOO 
Subject: PMC, IMC aligned. The ringdown PD/Lens removed. 

PMC and IMC were aligned on Friday (16th) and Today (19th).

The PD and lens for the ringdown experiment were removed as they were blocking the WFS.

ELOG V3.1.3-